• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name S. Sirca

    Log entry time 00:06:30 on November26,2008

    Entry number 250773

    keyword=Swing shift SUMMARY


    SWING shift summary: S. Sirca (SL), J.-C. Peng (TO), X. Ding (3rd)
    ==================================================================

    EVENTS
    ------
    16:00 took over from day shift, beam being restored
    19:30 (approx) start of beam tuning through Compton chicane
    21:40 MCC is ready to send CW
    21:50 Harp scans, see ELOG
    23:30 Problems with resetting Beamline, HV IOCs

    RUNS
    ----
    Rate scans with increasing raster, reference cell with 1bar N2, tune beam::
    #3917 raster 2x2 (MMC value)
    #3918 raster 3x4
    #3919 raster 4x5
    #3920 raster 5x6.6, observed rate increase here, see this and this HALOG entry.
    Reference cell runs:
    #3921 junk (wrong prescales)
    #3922 junk (some HV missing)
    #3923 5uA, empty reference cell